OpenStudy (anonymous):

(5/x-2) - (2/x+2) = (3/x^2-4)

OpenStudy (anonymous):

start with \[\frac{5(x+2)-2(x-2)}{x^2-4}=\frac{3}{x^2-4}\]

OpenStudy (anonymous):

\[5x+10-2x+4=3\] \[3x+14=3\] \[3x=-11\] \[x=-\frac{11}{3}\]

OpenStudy (anonymous):

so if i do another equation like this for example, (x/x+4) - (3/x-4) = (-22/2^2-16) how would i find the LCD ?

OpenStudy (anonymous):

same idea

OpenStudy (anonymous):

you have one denominator \[x+4\] one \[x-4\] and one \[x^2-16\] and as luck would have it \[(x+4)(x-4)=x^2-16\] so you can use \][x^2-16\] as your lcd

OpenStudy (anonymous):

\[x^2-16\] as lcd

OpenStudy (anonymous):

oh, that formula is (a-b)^2 right?